DIRC Operations Weekly Report

Period: Saturday 2006/07/01 - Friday 2006/07/07

Oncall: Jerome Bequilleux
Fast Monitoring: Rolf Andreassen, Jerome Bequilleux

File location: $BFROOT/www/Detector/DIRC/WeeklyReports/report_20060708.html
Plot directory: $BFROOT/www/Detector/DIRC/WeeklyReports/Plots/report_20060708

Contents


BaBar-PEP news

PEP peak lumi: ~11*1033 (previous week: 10.2*1033)

BaBar integrated lumi: 4.483 fb-1 (previous week: 3.98 fb-1)

Run 5b PEP delivery: 62.49 fb-1

Run 5 total delivery: 129.50 fb-1 Run 5 total BaBar recorded: 124.86 fb-1 including 9.82 fb-1 off-peak data

PEP activities this week:

  • Saturday 1st/Sunday 2nd: Abort kicker timing and vacuum problem(pump) both fixed. They had a single fill for 22 hours. However fill difficulties started on Sunday.
  • Wednesday 5th : -NFTR problem
                                   -lumi ratio set to 0.98;
                                   -UPS work is done and the battery is ok
  • Thursday 6th :PEP is trying to diagnose the HER aborts.

    PEP activities planned next week :

  • try to find the origin of the HER aborts that occurs arround 1200mA.

    Foreseen work opportunities (PEP down periods):

    BaBar activities/issues this week:

    DIRC activities

    See DIRC logbook for details

  • UPS work:
           -July 5th:Jerry disabled the SOB dump valve as there was maintenance work on the UPS system.

  • Degasing Pump:
           -July 4th: received 3 automatic emails from IR2 to warn that the degasing pump was off: 3 very close glitches in the readeout of the degasing pump variables. Nicolas added a new protection to prevent us from new future single glich alarms (See: single glitch plot).
           -July 7th: New alarm "degasing pump off", but not single glitches. Went to IR2 to see the hardware: everything was fine. Obviously, we have a readout problem; Nicolas changed the red alarm threshold from 15 to 60, so that we won't have the alarm again, for these conditions.

  • Instructions updates:
           -July 5th: the run taking while the board #3 of sector 4 (crate in slot 3) was off have been DQG-ed OK by Eli and by the PID expert (30 Jun06) ->next time, the faulty board wouldn't need to be replaced immediately. Information added in the ops. manual.
           -July 7th: Nicolas added long term instruction about the green alarms DRC:HV:UNDERVOLTAGE, OVERVOLTAGE, TRIP.

  • Worrying unexplained issue:
           -July 6th: During DAY shift, around 11:15am, ROMs crashes. In fact, all 12 DIRC front-end crates went off at the same time for about 20 seconds, creating the ROMs crash (reason remains unclear).
           -July 7th: We don't see very well the front-end crates fans having a speed going to 0 because of the switch off of front-end crates the day before-> due to the scale of plots(see Fans Monitoring Plots ); will be changed.

  • Investigating the need of changing hv crate in slot #3:
           -July 7th: We talked with Paul Stiles about the 3 failing boards we gave him coming from crate in slot #3. No apparent reasons to blame the crate -> we'll probably keep it untill the shutdown, unless new problems appear with it.

    Downtime due to the DIRC: 0 minutes.

    Backgrounds

    Background summary plot

    Scheduled activities

    Pending Work

    Manpower situation

    Links